Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page YouTube channel

BlueSky profile YouTube channel
P68-502

P68-502: Establishing the accuracy of asteroseismic mass, radius, and age estimates of giant stars using eclipsing binaries

P.I.: Brogaard

This is a Monitoring program using FIES.

Instrument setup

  • FIES

  • High-res fiber F4

Afternoon instructions

  • Run afternoon calibrations: fies-calibs -fib 4

Observing instructions

  • On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'

  • Go to directory: cd ~/obs/scripts/68-502/ and do ls -lrt

  • Start the requested OB generated script


Program Tracking
Proposal 68-502
Semester P68: 20231001 - 20240401
Principal Investigator Karsten Brogaard
Programme Establishing the accuracy of asteroseismic mass, radius, and age estimates of giant stars using eclipsing binaries
Instrument FIES
Time Allocation Multiple epochs of 7 Kepler and K2 targets. ~1 hr per obs.
Total Nights 39.0h
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 22:12 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
68-502 2023-10-04 00:00 Mon kic4663623 - The observation failed, because we did not see the target during fiber acquisition despite trying three times. The combination of calima and reflected moonlight was most likely the reason for the fail. Failed
68-502 2023-10-06 01:07 Mon kic9153621 FIGj060156 - FIGj060158 Observed with clear skies and variable seeing around 1.7'' Observed
68-502 2023-10-16 01:06 Mon epic211892898 FIGj160124 - FIGj160126 Last object of the night, observed during a clear sky. Observed
68-502 2023-10-18 02:13 Mon kic9153621, epic211892898 FIGj180097 - FIGj180099, FIGj180114 - FIGj180116 Clear skies, sub-arcsec seeing Observed
68-502 2023-10-22 00:00 Mon The target KIC4663623 was missed as the telescope was closed due to high humidity. Failed
68-502 2023-10-30 00:00 Mon Could not be observed due to weather Failed
68-502 2023-11-05 00:00 Mon Could not be observed because of weather. Failed
68-502 2023-11-07 01:07 Mon kic4663623 FIGk070099 - FIGk070101 Observed with clear skies but very high wind speeds. Dimm seeing above 1.5 arcsec. Observed
68-502 2023-11-10 01:06 Mon epic211892898 FIGk100086 - FIGk100088 There were thin clouds in some parts of the sky, but I did not see any overlapping the target. Observed
68-502 2023-11-15 01:07 Mon epic211892898 FIGk150118 - FIGk150120 Clear skies, high dust levels, variable seeing between 1 and 2 arcsec. Observed
68-502 2023-11-25 02:13 Mon kic9153621, epic211892898 FIGk250094 - FIGk250096, FIGk250117 - FIGk250119 Observed through thin clouds, with seeing around 1 arcsec. Observed
68-502 2023-11-30 00:00 Mon No observations due to weather. Failed
68-502 2023-12-01 01:06 Mon epic211503363_20231201 FIGl010074 - FIGl010076 Observed with clear skies, seeing around 1" Observed
68-502 2023-12-02 01:06 Mon epic211892898 FIGl020096 - FIGl020098 Observed with clear skies and variable seeing around 1.0-2.5 arcsec. Observed
68-502 2023-12-04 00:00 Mon Target couldn't be observed due to weather. Failed
68-502 2023-12-07 01:07 Mon epic248854690 FIGl070068 - FIGl070070 Seeing around 2 arcsec Observed
68-502 2023-12-22 00:00 Mon 100% humidity all night Failed
68-502 2023-12-24 00:00 Mon - - The monitoring target was missed as the telescope was closed to night humidity. Failed
68-502 2024-01-03 02:13 Mon epic248854690, epic212694013 FIHa030111 - FIHa030116 Observed
68-502 2024-01-05 01:07 Mon epic248847494 FIHa050088 - FIHa050090 Thin clouds and seeing around 1.0-1.5''. There have been problems with a rise in temperature/pressure of the CCD, which made it more difficult to check the fiber guiding. The target seemed centered on the fiber and the exposure meter was counting, but the guiding is not certain. Observed
68-502 2024-01-06 01:07 Mon epic211892898 FIHa060068 - FIHa060070 Clear skies and 1'' seeing. Observed
68-502 2024-01-07 01:09 Mon epic211892898 FIHa070063 - FIHa070066 Observed with thin clouds. Seeing was around 1 arcsec. S/N estimates from exposure meter were 45 (O9), 48 (F3), 62 (G7) and 86 (M2). Observed
68-502 2024-01-13 01:12 Mon epic211892898 FIHa130094 - FIHa130096 Quite cloudy but guiding never lost and S/N of approx 40 obtained according to exposure meter. Observed
68-502 2024-01-24 00:00 Mon - - The target was missed due to bad weather Failed
68-502 2024-01-27 00:00 Mon Could not be observed due to moon distance. Failed
68-502 2024-01-28 01:08 Mon epic212694013 FIHa280098 - FIHa280100 Clear skies with some dust (around 35 micrograms per cubic meter or less). Seeing below 1 arcsec. The Moon was relatively close, so the Moon phase limit was not fulfilled. Vignetting occurred for the first 10 minutes or less as the lower hatch was kept closed throughout the night. Afterwards, the target had risen above the vignetting limit. Observed
68-502 2024-02-01 00:00 Mon Could not observe any of the targets due to a technical issue with stancam (FR4410) Failed
68-502 2024-02-15 00:58 Mon epic248847494 FIHb150070 - FIHb150072 Observed with clear skies and seeing of approx. 0.6 arcseconds Observed
68-502 2024-03-04 00:00 Mon epic248847494 FIHc040063 - FIHc040065 Good night. Observed
68-502 2024-03-29 00:00 Mon Unable to observe target due to high humidity and freezing conditions all night. Failed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: September 27 2023